OpenStudy (anonymous):

The price of bronze has increased by 10% per year from 2000. In the year 2000, Harry bought a bronze medal for $120. Which of the following functions f(x) can be used to represent the price of the medal x years after 2000?

OpenStudy (anonymous):

(A) f(x) = 120(0.9)^x (B) f(x) = 10(120)^x (C) f(x) = 120(1.1)^x (D) f(x) = 10(1.1)^x

OpenStudy (anonymous):

@Abhisar @cwrw238 @SolomonZelman @TuringTest @e.mccormick I need help! Please someone help me!!

OpenStudy (anonymous):

@mathmath333 @Michele_Laino @iGreen

OpenStudy (anonymous):

\[A=P(1+r)^{t}\] This is your equation for exponential growth. where A= new amount, P= starting amount, r= rate, and t=time

OpenStudy (anonymous):

knowing this equation, plug in the values corresponding to the problem and this should give you your answer.

OpenStudy (anonymous):

Okay, so that would make f(x) = A 120 = P

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

when you add 1 to the 0.1 that would make it equal to 1.1

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

and t = time, meaning that x = t

OpenStudy (anonymous):

you've got it

OpenStudy (anonymous):

So then the answer is C.

OpenStudy (anonymous):

That's it.

OpenStudy (anonymous):

Thank YOU SO MUCH! :)

OpenStudy (anonymous):

No problem.